OpenStudy (anonymous):

which of the following equations is of a parabola with a vertex at (0 -3)

OpenStudy (anonymous):

can u help me @LizzyLove<3

OpenStudy (phi):

do you know the formula for a parabola ?

OpenStudy (anonymous):

not really

OpenStudy (phi):

the most convenient form is the equation of a parabola in vertex form.

OpenStudy (anonymous):

Maybe phi has it all under control:) @Shannon20150

OpenStudy (anonymous):

idk that form

OpenStudy (phi):

See http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php for the two forms

OpenStudy (phi):

If this is multiple choice what are your choices?

OpenStudy (phi):

But if you look at the site, the vertex form is y = a(x-h)^2+k you are told (h,k) is (0,-3) if we put those numbers in for h and k in the standard form we get y= a(x-0)^2+ -3 or y= a x^2 -3 so if you see anything like this as your choice (where a can be any number), that is the answer

OpenStudy (anonymous):

Y=(x-3)^2 Y=(x+3)^2 Y=x^2-3 Y=x^2+3 I think that it is the 1st one

OpenStudy (phi):

maybe you should read by previous post more carefully

OpenStudy (phi):

if you can't remember all this, you can try this. you know (0 , -3) is on the parabola. make x =0, and see if any of the choices gives you y= -3 this is not the best way... but it might work in a pinch. (they could give a parabola where 0,-3 is on the parabola, but not the vertex, in which case you are in trouble.

OpenStudy (anonymous):

so how would I set it up

OpenStudy (anonymous):

@some_someone

OpenStudy (phi):

what part are you not getting ? y = a(x-h)^2+k you are told (h,k) is (0,-3) if we put those numbers in for h and k in the standard form we get y= a(x-0)^2+ -3 or y= a x^2 -3

OpenStudy (anonymous):

ohhhhhhhhhh okay

OpenStudy (phi):

what is the answer ? (remember a is any number) match y= a x^2 -3 to your choices